Find the value of the acceleration due to gravity of a body of mass 0.6kg and at a distance of 6×10-6 from the moon [G=6.7×10-11Nm2kg-2]

1
Expert's answer
2021-01-22T11:29:04-0500
"g_M=\dfrac{GM_M}{(R_{M}+h)^2},""g_M=\dfrac{6.7\cdot10^{-11}\ \dfrac{Nm^2}{kg^2}\cdot 7.36\cdot10^{22}\ kg}{(1.737\cdot10^6\ m+6\cdot10^6\ m)^2}=0.082\ \dfrac{m}{s^2}."

Answer:

"g_M=0.082\ \dfrac{m}{s^2}"